# TEXasGAN: Tactile Texture Exploration and Synthesis System Using Generative Adversarial Network

### v1

@@ I Introduction
There have been many studies concentrated on the reconstruction of vibration signals through the utilization of spectrograms derived from sensor data capturing physical surface interactions, and also gave the optimized spectrgrams that match the desired texture pattern well using generative models [4, 7]. In pursuit of a broader and more versatile spectrogram of vibrotactile signals, deep learning based generative models, particularly Generative Adversatial Network (GAN), have been enlisted. These methods can effectively optimize the objective function and yield satisfactory generative results. Building upon this, we also recognize that human expert knowledge may be another crucial factor influencing the optimization, particularly in tasks related to subjective perception.
@@
In this paper, we introduce TEXasGAN, a system designed to allow users to create customized vibrotactile feedback based on their subjective perceptions, as shown in Fig. 1. Our approach utilizes the Differential Subspace Search (DSS) method [9], enabling users to explore a high-dimensional latent space through a 1-D slider. To address the challenge of reaching every point in the latent space, we also implemented an initialization method to provide a good starting point. We trained GAN-based neural network models and used the pre-trained generator to create spectrograms from latent vectors optimized via DSS. The user interface features a simple 1-D slider, enabling users to experience and evaluate generated tactile vibrations. Through iterative exploration and selection, users can identify the optimal slider position that yields vibrations closest to their desired stimulus. Our two-stage user experiment, consisting of generation and evaluation phases, demonstrated that TEXasGAN can produce distinguishable vibrations. Despite the extensive research on human-in-the-loop approaches in various interaction methods, their application in haptics remains underexplored. We hope our work contributes to the development of more intuitive and realistic haptic feedback systems.
@@ II-A Vibration Tactile Presentation
Significant progress has been achieved in the advancement of vibrotactile haptic devices, making them a central focus of investigation within the domain of tactile presentation systems. Numerous studies have been dedicated to the replication of diverse sensory experiences. For instance, through the utilization of piezo actuators, vibrotactile feedback on touchscreens has been employed to simulate the tactile sensation of physical buttons, as exemplified in the work by Sadia et al. [10]. Beyond touchscreen interactions, wearable devices like FingerX [11] have emerged to offer users a heightened sense of object shapes within VR environments.
@@ II-B Generative Neural Network
Autoencoders, owing to their relatively straightforward architecture and unsupervised training paradigm, have been employed in some studies to serve as tools for compressing and reconstructing tactile signals [16, 17]. Nevertheless, the compression process can introduce oversmoothing, resulting in the loss of fine-grained details, making autoencoders difficult to reproduce detailed texture related vibration signals. Although increasing the dimension of the compressed latent vector can solve the problem, it also makes operations like optimization more difficult. Therefore, the use of generative models, rather than a decoder within the autoencoder framework, to reconstruct vibration signals has become a prudent choice.
GAN, a prominent generative neural network model, has garnered widespread recognition for its applications in image generation and is progressively finding utility in diverse tactile reconstruction tasks as well [7, 4, 18]. GAN primarily comprises two components: the generator and the discriminator. The generator is tasked with generating samples from a specified distribution, commonly Gaussian noise, while the discriminator’s role is to differentiate between samples generated by the generator and real samples. Through an iterative training process, the discriminator’s proficiency in discerning genuine from generated samples continually improves. Simultaneously, guided by the discriminator, the generator becomes increasingly good at producing more real samples.
@@ III-B Generator Design
To ensure the effectiveness of the generator, it is essential to employ an appropriate discriminator. A discriminator that performs exceptionally well or very poorly can hinder the training of the generator. In our study, we utilized a discriminator with a structure akin to the one presented in [22]. This discriminator comprises 8 convolutional layers, albeit with adjustments to the channel numbers to align it with the specific requirements of our task. The comprehensive architecture of the Generative Adversarial Network (GAN) is depicted in Fig. 3.
@@ III-C Generation with Conditional Knowledge
We initialized the encoder using a pre-trained ResNet-50 model that was trained on the ImageNet1K_V2 dataset, and introduced an FC layer to resize the spectrogram to the shape corresponding to the pre-trained model input. The generator and the spectrogram discriminator follow the network structures illustrated in Fig. 3, maintaining consistency with those architectures. The latent discriminator contains 2 FC layers, and 1 output layer for the discrimination and 1 output layer for the classification. To apply the label as conditional information, we used one-hot labels and used an FC layer to resize the label vector to the dimension corresponding to the feature map. And then the label information will be added to the feature map as a channel.
@@ IV-A Data Preparation
Seven groups (G1–G4, G6, G8, and G9) were selected from the total of nine groups in LMT108. The G5 and G7 groups were excluded from the study, as the original stimuli were found to evoke a tactile sensation that was similar to that of other groups. And we selected 2 kinds of textures from each groups and built a training dataset containing 14 classes.
@@
The original dataset contains about 5 seconds sound recordings of different surfaces. Here we used a sliding window of 48 $\times$ 320 corresponding to about 1.5 seconds, and moved it on transformed spectrograms, so that we can obtain segmentations from different beginning time. The moving step length was set at 5 units (about 23 milliseconds), and produced a total of 24480 spectrogram segments.
@@ IV-C Initialization
To bring $z$ closer to the target area, the principle of our search strategy is: Choose near but different new $z$ for “So-so”, and far new $z$ for “Bad”. The average distance of vectors in the latent space $dis_{avg}$ was calculated. Here we define a parameter $step=\frac{1}{8}\times dis_{avg}$ as the unit of the moving distance during the exploration. Inspired by the tabu search, visited latent vectors are added to the tabu list. In this study, a specified radius $r$ is enforced for vectors within the tabu list (here $r$ is defined as the average distance $step$). Consequently, vectors within this range are prohibited, allowing for the exploration of larger regions. In this study, the length of the tabu list is set to 5, meaning that the areas surrounding the 5 most recently visited latent vectors are prohibited from the further exploration. Following the search principle mentioned, when the user selects “So-so”, we choose a new $z$ within the range of (0.5 $\times$ $step$, 2 $\times$ $step$) from the current $z$. If the user chooses “Bad”, we select a new $z$ from distances beyond 2 $\times$ $step$ from the current $z$.
@@ IV-E Stage I: Generation
To assess the generative performance of our TEXasGAN, we designed a two-stage experiment that includes a generation experiment and an evaluation experiment. To validate whether TEXasGAN can generate new samples different from the training data, we selected 5 classes from LMT108 that differ from the training set as generation targets, as shown in Fig. 6.
Due to the free movement of the collection device on surfaces during the vibration acquisition process [28], vibration is not uniformly distributed on the time axis. This implies that not all samples can effectively represent the characteristics of a particular class. To provide subjects with more reliable target vibrations, facilitating better execution of the generation task, we manually selected 3 relatively uniform and representative vibrations for each class. This resulted in a total of 15 target vibrations, and subjects were tasked with generating samples as similar as possible to these vibrations.
@@
In order to cover as many generated results as possible in the subsequent pairing task of the evaluation experiment without overwhelming the subjects, we limited the number of generated samples. In the first stage, we invited 3 subjects (3 males, aged 21-26), who totally generated 45 samples (9 for each class). These three subjects will not be involved in the evaluation phase to eliminate the impact of memory.
@@ IV-F Stage II: Evaluation
We designed a classification task to validate the similarity between generated vibrations and real surface vibrations. The classification task comprised three parts:
1. Select the correct class for the real vibration (the control group): In this part, all 15 real vibrations were presented in a shuffled order to the subject. Each time, one vibration was shown, and 5 other randomly selected real vibrations from the 5 classes were also presented (excluding the vibration to be classified). The subject was required to choose the vibration that most closely resembled the presented vibration. This process was repeated three times, resulting in a total of 45 classifications. The aim of this part was to validate the premise of this study, which is that people can perceive differences between vibration stimuli, understand their characteristics, and correctly classify them into corresponding categories. Simultaneously, the conclusions from this control group were used to verify the fidelity of the generated results. We also asked subjects to describe the features of these 5 kinds of vibrations and conduct the following tasks according to these features.
2. Select the correct class for the generated vibration. The procedure of the second part was similar to the first part. However, instead of the real vibration, the vibration to be classified was replaced by the generated vibration. The 45 generated vibrations were shuffled and presented to the subject. The subject was required to select the vibration that most closely resembled the vibration generated from the five options, each of the five categories. This process was only conducted once to maintain an equal task number with the control group. We asked subjects to describe the vibration features and compare with the control group.
3. Select the corresponding target vibration for the vibration generated. The third part aimed to validate if subjects can identify differences between samples from a same class. The 45 shuffled generated vibrations were shown to the subjects. However, the references were changed to three vibrations in the same corresponding class. The subject needed to select the correct target vibration corresponding to the given generated vibration. The total number of classification tasks in this part was also 45.
Twenty subjects (11 males, 9 females, aged 22-27) took part in this stage. The entire process of the evaluation experiment lasted about 1.5 hours, including the tutorial and rest breaks.
@@ V Results
In this section we will first present the performance of the trained neural network model, and then present the results of the 2-stage user experiments.
@@ V-B Problems in Other Alternatives
At first we experimented with models based on the autoencoder architecture, namely autoencoder, VAE, and AAE. These models were easy to train but suffered from severe over-smoothing issues. As we needed feature vectors suitable for optimization, we aimed for a relatively small feature space dimension, such as the 128 dimensions used in this study. However, 128 proved insufficient for the original data, resulting in significant information loss when encoded into this latent space. Consequently, the generated spectrograms lacked detailed information, leading to substantial differences between the reconstructed vibrations and the original ones.
We also implemented GAN within our framework; however, it tended to suffer from mode collapse. Although it could generate spectrograms with more details compared to autoencoders, the model exhibited a strong tendency to produce similar results for different input. This behavior rendered it unsuitable for our optimization based on input vectors. In addition, GAN gives results from random noise generation, which brings about a mixed latent space. Therefore, as mentioned earlier in Sec. III-C and Sec. IV-C, the generated vibrations underwent drastic changes when the user continuously moved the slider. Unlike the original DSS research, which originally focused on visual stimuli, we emphasize tactile tasks, which have a much lower sensitivity compared to visual stimuli. Therefore, the abrupt changes pose difficulties for users in touching and comparing vibrations, making it nearly impossible for users to select an appropriate slider position.
@@ V-C User Study
However, spectrograms alone cannot fully capture the subjective experiences of the subjects. Therefore, we mainly focus on analyzing the results of Stage II, the classification task. From Fig. 9 we can see an overview of the result of the classification task. In the five-category task, subjects achieved an accuracy of 55.9% when judging the generated vibrations, compared to about 72.8% for the real vibrations. The control group results indicate that subjects can distinguish between different surface vibration stimuli. Fig. 9 shows that subjects in the generation group and the control group exhibited similar accuracy trends for these 5 kinds of different vibrations. This preliminary evidence suggests that our TEXasGAN can generate distinct vibration patterns that subjects can differentiate.
@@
We constructed a linear regression model with interaction terms, using control group accuracy, subject ID, and vibration class as independent variables, and the generated group accuracy as the response. The model can be expressed as:
$acc_{G}\sim 1+c*i+c*acc_{R}+i*acc_{R}.$ (6)
In the equation, $acc_{G}$ represents the generated group accuracy, $c$ represents the class, $i$ represents the subject ID and $acc_{R}$ represents the control group accuracy. Before the fit process, we removed outliers that were more than two scaled median absolute deviations from the median, obtained 93 sets of accuracy. Then we performed a stepwise regression to select significant variables. In this way, we can obtain a model like:
$acc_{G}\sim 1+c*acc_{R}.$ (7)
This model keeps category $c$ ($p=0.0699$), control group accuracy $acc_{R}$ ($p=1.1785\times 10^{-8}$), and their interaction term $c*acc_{R}$ ($p=0.0012$). This shows that there is indeed a certain relationship between $acc_{R}$ and $acc_{G}$. At the same time, the subject $i$ ($p=0.4739$) is not a significant variable, indicating that this is a relatively common rule among subjects. The model $R^{2}=0.423$, means there is some certain relationship; $F=21.7$ with $p=1.19\times 10^{-10}$, means that the model is credited.
@@
Figure 13: The in-class accuracy of each class.

### v2

@@ I Introduction
There have been many studies concentrated on the reconstruction of vibration signals through the utilization of spectrograms derived from sensor data capturing physical surface interactions, and also gave the optimized spectrograms that match the desired texture pattern well using generative models [4, 7]. In pursuit of a broader and more versatile spectrogram of vibrotactile signals, deep learning based generative models, particularly Generative Adversatial Network (GAN), have been enlisted. These methods can effectively optimize the objective function and yield satisfactory generative results. Building upon this, we also recognize that human expert knowledge may be another crucial factor influencing the optimization, particularly in tasks related to subjective perception.
@@
In this paper, we introduce TEXasGAN, a user-friendly system integrating GAN and DSS for user-driven optimization of vibrotactile stimuli, as shown in Fig. 1. A simplified 1-D slider interface was developed for easy exploration of high-dimensional tactile latent space. We also emphasized the role of human perception in optimizing and evaluating tactile feedback, highlighting the significance of integrating subjective preferences into the design process. The experiment results show that subjects achieved a 58% accuracy rate in distinguishing the categories of generated vibrations in a five-class task, compared to 74% for real vibrations. We also used a linear regression model to demonstrate a significant relationship between these two groups of accuracy rates, indicating our system’s capability to recreate distinct vibration characteristics based on user preferences.
We have introduced a novel idea for optimizing tactile parameters, enabling control over high-dimensional feature spaces. By integrating various network architectures, we have created a generative model tailored specifically for optimizing tactile parameters. Despite the extensive research on human-in-the-loop approaches in various interaction methods, their application in haptics remains underexplored. We are also trying to bring We hope our work contributes to the development of more intuitive and realistic haptic feedback systems.
@@ II-A Vibration Tactile Presentation
Significant progress in vibrotactile haptic devices has made them a key focus in tactile presentation systems. Many studies aim to replicate various sensory experiences. For example, piezo actuators are used to simulate physical button sensations on touchscreens, as demonstrated by Sadia et al. [10]. Additionally, wearable devices like FingerX [11] enhance the sense of object shapes in VR environments.
@@ II-B Generative Neural Network
Autoencoders, due to their straightforward architecture and unsupervised training approach, have been used in some studies for compressing and reconstructing tactile signals [16, 17]. Nevertheless, the compression process can introduce oversmoothing, resulting in the loss of fine-grained details, making autoencoders difficult to reproduce detailed texture related vibration signals. Although increasing the dimension of the compressed latent vector can solve the problem, it also makes operations like optimization more difficult. Therefore, the use of generative models, rather than a decoder within the autoencoder framework, to reconstruct vibration signals has become a prudent choice.
Generative Adversarial Networks (GANs), a notable generative model, are widely recognized for their success in image generation and are increasingly applied to tactile reconstruction tasks [7, 4, 18]. A GAN consists of two main components: the generator and the discriminator. The generator produces samples from a specified distribution, often Gaussian noise, while the discriminator’s role is to distinguish between generated samples and real ones. Through iterative training, the discriminator’s ability to differentiate between genuine and generated samples improves, which in turn enhances the generator’s capability to produce increasingly realistic samples.
@@ III-B Generator Design
To ensure the effectiveness of the generator, it is essential to employ an appropriate discriminator. A discriminator that performs exceptionally well or very poorly can hinder the training of the generator. In our study, we utilized a discriminator with a structure similar to the one presented in [22]. This discriminator comprises eight convolutional layers, albeit with adjustments to the channel numbers to align it with the specific requirements of our task. The comprehensive architecture of the Generative Adversarial Network (GAN) is depicted in Fig. 3.
@@ III-C Generation with Conditional Knowledge
We initialized the encoder using a pre-trained ResNet-50 model that was trained on the ImageNet1K_V2 dataset, and introduced an FC layer to resize the spectrogram to the shape corresponding to the pre-trained model input. The generator and the spectrogram discriminator follow the network structures illustrated in Fig. 3, maintaining consistency with those architectures. The latent discriminator contains two FC layers, and 1 output layer for the discrimination and one output layer for the classification. To apply the label as conditional information, we used one-hot labels and used an FC layer to resize the label vector to the dimension corresponding to the feature map. And then the label information will be added to the feature map as a channel.
@@ IV-A Data Preparation
Seven groups (G1–G4, G6, G8, and G9) were selected from the total of nine groups in LMT108. The G5 and G7 groups were excluded from the study, as the original stimuli were found to evoke a tactile sensation that was similar to that of other groups. And we selected two kinds of textures from each groups and built a training dataset containing fourteen classes.
@@
The original dataset contains about five seconds sound recordings of different surfaces. Here we used a sliding window of 48 $\times$ 320 corresponding to about 1.5 seconds, and moved it on transformed spectrograms, so that we can obtain segmentations from different beginning time. The moving step length was set at five units (about 23 milliseconds), and produced a total of 24480 spectrogram segments.
@@ IV-C Initialization
To bring $z$ closer to the target area, the principle of our search strategy is: Choose near but different new $z$ for “So-so”, and far new $z$ for “Bad”. The average distance of vectors in the latent space $dis_{avg}$ was calculated. Here we define a parameter $step=\frac{1}{8}\times dis_{avg}$ as the unit of the moving distance during the exploration. Inspired by the tabu search, visited latent vectors are added to the tabu list. In this study, a specified radius $r$ is enforced for vectors within the tabu list (here $r$ is defined as the average distance $step$). Consequently, vectors within this range are prohibited, allowing for the exploration of larger regions. In this study, the length of the tabu list is set to five, meaning that the areas surrounding the five most recently visited latent vectors are prohibited from the further exploration. Following the search principle mentioned, when the user selects “So-so”, we choose a new $z$ within the range of (0.5 $\times$ $step$, 2 $\times$ $step$) from the current $z$. If the user chooses “Bad”, we select a new $z$ from distances beyond 2 $\times$ $step$ from the current $z$.
@@ IV-E Stage I: Generation
To assess the generative performance of our TEXasGAN, we designed a two-stage experiment that includes a generation experiment and an evaluation experiment. To validate whether TEXasGAN can generate new samples different from the training data, we selected five classes from LMT108 that differ from the training set as generation targets, as shown in Fig. 6.
Due to the free movement of the collection device on surfaces during the vibration acquisition process [28], vibration is not uniformly distributed on the time axis. This implies that not all samples can effectively represent the characteristics of a particular class. To provide subjects with more reliable target vibrations, facilitating better execution of the generation task, we manually selected three relatively uniform and representative vibrations for each class. This resulted in a total of fifteen target vibrations, and subjects were tasked with generating samples as similar as possible to these vibrations.
@@
In order to cover as many generated results as possible in the subsequent pairing task of the evaluation experiment without overwhelming the subjects, we limited the number of generated samples. In the first stage, we invited three subjects (three males, aged 21-26), who totally generated forty-five samples (nine for each class). These three subjects will not be involved in the evaluation phase to eliminate the impact of memory.
Figure 9: The user interface of the Stage II. Subjects can click the upper button to play the vibration to classify. Then subjects should select a best matching one from the below real vibration samples.
@@ IV-F Stage II: Evaluation
We designed a classification task to validate the similarity between the vibrations generated and the real vibrations of the surface. The user interface for this part is shown in Fig. 9. The classification task comprised three parts:
1. Select the correct class for the real vibration (the control group): In this part, all fifteen real vibrations were presented in a shuffled order to the subject. For each trial, one real vibration was selected for classification, and subjects could click a button to play this vibration. Also, five real vibrations from each class were presented. These were randomly selected from each class, excluding the vibration to be classified shown above. The indices (1, 2, 3, 4, 5) corresponding to the vibration classes (G2, G3, G4, G6, G8). Subjects could play these vibrations and were required to choose the class that most closely resembled the one being classified. This process was repeated three times, resulting in a total of forty-five classifications. This part aimed to validate the study’s premise that people can perceive differences between vibration stimuli, understand their characteristics, and accurately classify them. The results from this control group were used to assess the accuracy of the generated vibrations. Subjects were also asked to describe the features of the five types of vibrations and perform subsequent tasks based on these descriptions.
2. Select the correct class for the generated vibration. The procedure for this part was similar to the first, but focused on classifying generated vibrations. The forty-five generated vibrations were shuffled and presented to the subjects for classification. The vibrations below were still randomly selected from the real vibrations of each class. Subjects were required to select the class that most closely resembled the vibration generated from these five options. This process was only conducted once to maintain an equal task number with the control group. We asked subjects to describe the vibration features and compare with the control group.
3. Select the corresponding target vibration for the vibration generated. The third part aimed to validate whether subjects can identify differences between samples in the same class. The forty-five shuffled generated vibrations were shown to the subjects, same as in the second part above. However, instead of five options, subjects were presented with three options, all belonging to the same class as the displayed generated vibration. The reason is, for each class, there were three real vibrations serving as targets. Therefore, subjects needed to select the correct target vibration from the below three options corresponding to the given generated vibration. The total number of classification tasks in this part was also forty-five, decided by the number of generated vibrations.
Twenty subjects (eleven males, nine females, aged 22-27) took part in this stage. The entire process of the evaluation experiment lasted about 1.5 hours, including the tutorial and rest breaks.
@@ V Results
In this section, we will first present the performance of the trained neural network model, and then present the results of the two-stage user experiments.
@@ V-B Problems in Other Alternatives
Initially, we experimented with models based on the autoencoder architecture, including autoencoders, VAEs, and AAEs. While these models were relatively easy to train, they suffered from severe over-smoothing issues. Given our need for feature vectors suitable for optimization, we targeted a relatively small feature space dimension, such as the 128 dimensions used in this study. However, this dimensionality proved inadequate for the original data, resulting in significant information loss when encoded into this latent space. Consequently, the generated spectrograms lacked detailed information, leading to noticeable differences between the reconstructed vibrations and the originals.
We also incorporated GANs into our framework. Although GANs could generate spectrograms with more detail compared to autoencoders, they were prone to mode collapse, producing similar results for different inputs. This issue rendered GANs unsuitable for our optimization approach based on input vectors. Additionally, GANs generate results from random noise, resulting in a mixed latent space. As a result, the generated vibrations underwent drastic changes when users moved the slider continuously. Unlike the original DSS research, which focused on visual stimuli, our work emphasizes tactile tasks, which are less sensitive than visual stimuli. Consequently, these abrupt changes made it difficult for users to touch and compare vibrations, complicating the task of selecting an appropriate slider position.
@@ V-C User Study
However, spectrograms alone cannot fully capture the subjective experiences of the subjects. Therefore, we mainly focus on analyzing the results of Stage II, the classification task. From Fig. 10 we can see an overview of the result of the classification task. In the five-category task, subjects achieved an accuracy of 55.9% when judging the generated vibrations (experimental group, experiment 2), compared to about 72.8% for the real vibrations (control group, experiment 1). The control group results indicate that subjects can distinguish between different surface vibration stimuli. Fig. 10 shows that subjects in the generation group and the control group exhibited similar accuracy trends for these 5 kinds of different vibrations. This preliminary evidence suggests that our TEXasGAN can generate distinct vibration patterns that subjects can differentiate.
@@
We constructed a linear regression model with interaction terms, using control group accuracy, subject ID, and vibration class as independent variables, and the generated group accuracy as the response. Before the fit process, we removed outliers that were more than two scaled median absolute deviations from the median, obtained 93 sets of accuracy. The regressed model can be expressed as:
$acc_{G}=0.0349*c-0.0017*i+0.9490*acc_{R}+$ (6)
$0.0011*c*i-0.1983*c*acc_{R}-0.0062*i*acc_{R}.$
In the equation, $acc_{G}$ represents the generated group accuracy, $c$ represents the class ID, $i$ represents the subject ID and $acc_{R}$ represents the control group accuracy. The $p$-values of each term are shown in Table I. The analysis reveals that the accuracy of the control group, $acc_{R}$, has the most significant impact, along with the interaction term between class $c$ and $acc_{R}$, both of which are statistically significant. This suggests a meaningful relationship between $acc_{G}$ and $acc_{R}$, indicating that our model successfully reconstructs recognizable features of real target samples. The significance of the interaction term $c*acc_{R}$ implies that the model’s generation capability varies across different classes. Additionally, terms involving the subject ID $i$ are not significant, suggesting that the observed patterns are consistent across subjects.
TABLE I: Coefficients and $p$-values of each factor in the model.
Coefficient $p$-value
$c$ 0.0349 0.5236
$i$0.0017 0.8963
$acc_{R}$ 0.9490 8.4611×$\times 10^{-5}$
$c*i$ 0.0011 0.6998
$c*acc_{R}$0.1983 0.0053
$i*acc_{R}$0.0062 0.7012
Then we performed a stepwise regression to select some variables. We can obtain a model like:
$acc_{G}=0.10245*c+1.0914*acc_{R}-0.2657*c*acc_{R}.$ (7)
This model keeps class $c$ ($p=0.0699$), control group accuracy $acc_{R}$ ($p=1.1785\times 10^{-8}$), and their interaction term $c*acc_{R}$ ($p=0.0012$). The $acc_{R}$ is still the most dominant term. The model still shows the difference in generation ability between classes, even larger than the previous model. This is reasonable according to our experiment results. The model $R^{2}=0.423$, means there is some certain correlation; $F=21.7$ with $p=1.19\times 10^{-10}$, means that the model is credited.
@@
Figure 14: The in-class accuracy of each class. The accuracy of each class is just over 33%, which means that although there are differences between the samples, they are difficult to distinguish.